Is it possible to find a spherical triangle having the following sides? Explain why?

a = 120° b = 125° c = 130°
1
Expert's answer
2020-11-25T17:59:16-0500

yes , it is possible

by the law of cosines of sides and

law of sines



now ,

using law of cosines for sides :

cos a = cos b cos c + sin b sin c cosA

cos 120"\degree" = cos 125"\degree"cos 130"\degree" + sin 125"\degree" sin 130"\degree" cosA

A = 77"\degree"93'

using law of sines:

"\frac{sinC}{sinc}" = "\frac{sinA}{sina}"


"\frac{sinC}{sin130\degree}" = "\frac{sin77\degree93'}{sin120\degree}"


C = 60"\degree"48'


"\frac{sinB}{sinb}" = "\frac{sinA}{sina}"


"\frac{sinB}{sin125\degree}" = "\frac{sin77\degree93'}{sin120\degree}"


B = 68"\degree"51'


now we also can see spherical excess

E = (A+B+C) - 180"\degree"

E = (77"\degree"93' + 68"\degree"51' + 60"\degree"48' )-180"\degree"

E = 26"\degree"92'